OpenStudy (anonymous):

I need some help with graphing quadratics. How do I find the x and y-intercepts in both y=(x-h)^2+k and y=ax^2+bx+c forms?

OpenStudy (anonymous):

for x intercept, assume y=0 and for y intercept, assume x= 0........substitute these value in equation and you will get answer

OpenStudy (saifoo.khan):

y=ax^2+bx+c solve the quadratic equation to get the roots (x-intercepts) when y intercept put x = 0 and solve.

OpenStudy (saifoo.khan):

y=(x-h)^2+k this form is basically used to find thevertex.

OpenStudy (saifoo.khan):

the vertex*

OpenStudy (anonymous):

y=(x-1)^2-16 By substituting x for zero, I got -15. So the y-intercept is 0,-15. I'm still confused on how to find the x-intercept.

OpenStudy (saifoo.khan):

correct. now, y=(x-1)^2-16 (x-1)(x-1) - 16 = 0 solve this using foil.

OpenStudy (anonymous):

x^2-2x-15

OpenStudy (saifoo.khan):

correct. now use the factorization or use the quadratic formula to solve this quadratic equation.

OpenStudy (anonymous):

3,0 and -5,0

OpenStudy (saifoo.khan):

wouldn't it be, -3,0 and 5,0

OpenStudy (anonymous):

No, because when you add -3 and 5 you get 2, not -2.

OpenStudy (saifoo.khan):

x^2-2x-15 = x^2 - 5x + 3x - 15 = 0 x(x-5) + 3(x-5) = 0 (x+3)(x-5) = 0 (x+3) = 0 | (x+5) = 0 x = -3 | x = -5